OpenStudy (anonymous):

find the measures of two angles, one positive and one negative, which are the co terminals with the given angle. 2pi/3

OpenStudy (anonymous):

\[\frac{2\pi }{ 3}\]

OpenStudy (anonymous):

my question is do i convert it to 120 degree

OpenStudy (anonymous):

360-120=240 and 360+120=480 is this correct?